OpenStudy (anonymous):

Write an equation of the sine function with: amplitude = 7 period = 3π phase shift = π vertical shift = -7

OpenStudy (anonymous):

So lets think about what these things mean... Are you familiar with amplitude?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

OK, Amplitude is how much to 'amplify' the function by.

OpenStudy (anonymous):

or you could say, multiply.

OpenStudy (anonymous):

If you have 5Sin(x) then its amplitude is 5 (the wave will hit a maximum of 5 and a minimum of -5.

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

So hopefully that explains amplitude. The period is the length of a wave before it starts repeating itself. The usual period of a sine function is 2Pi as the wave will begin to repeat after that point in time. I.e. period of Sin(x) is 2Pi. Is this making sense so far?

OpenStudy (anonymous):

so its 7sinx

OpenStudy (anonymous):

yep, thats right, 7 sin x (for the amplitude part)

OpenStudy (unklerhaukus):

\[y(x)=A\sin(\omega t+\phi)+b\] \(A\) is the Amplitude \(\omega\) is the angular frequency \(\frac{\omega}{2\pi}=\)frequency of oscillation \(\tau\) is the period and \(\tau=\frac{2\pi}{\omega}\) \(\phi\) is the phase shift \(b\) is the vertical shift

OpenStudy (anonymous):

You are best of taking a look at http://www.intmath.com/trigonometric-graphs/trigo-graph-intro.php. This will explain it properly.